consumer has money income of USD 400 and consumes x and y with price of x and y are equals 2 and 5 respectively . U=X^1/3 Y^2/3

A) Calculate the demand function for x and y

B) calculate the utility maximize levels of x and y

C) compute the maximize utility that consumer drive from consumption of x and y


1
Expert's answer
2022-01-29T11:06:26-0500

Solution:

A.). Demand function for X:

"\frac{MUx}{MUy} = \frac{Px}{Py}"


"\frac{Y}{Px} = \frac{X}{Py}"


Y = "\frac{Px}{Py}X"

M = PxX + PyY

M = PxX + Py("\frac{Px}{Py}X")

M = PxX + PxX

M = 2PxX

X = "\frac{M}{2Px}"

 

Demand function for Y:

"\frac{MUx}{MUy} = \frac{Px}{Py}"


"\frac{Y}{Px} = \frac{X}{Py}"

X = "\frac{Py}{Px}Y"

M = PxX + PyY

M = Px("\frac{Py}{Px}Y") +PxX

M = PyY + PyY

M = 2PyY

Y = "\frac{M}{2Py}"

 

B.). Derive MRSXY = "\frac{MUx}{MUy} = \frac{Px}{Py}"

U= X1/3Y2/3

MUx = "\frac{\partial U} {\partial X} = \frac{1} {3}X^{-\frac{2}{3} } Y ^{\frac{2}{3}}"


MUy = "\frac{\partial U} {\partial X} = \frac{2} {3}X^{\frac{1}{3} } Y ^{-\frac{1}{3}}"

 

Set "\frac{MUx}{MUy} = \frac{Px}{Py}"


"\frac{1} {3}"X-2/3Y2/3 "\div" "\frac{2} {3}"X1/3Y-1/3= "\frac{2} {5}"


X = 1.23Y

Budget constraint: M = PxX + PyY

400 = 2X + 5Y

Substitute the value of X in budget constraint to derive Y:

400 = 2(1.23Y) + 5Y = 2.46Y + 5Y = 7.46Y

Y = 53.6

X = 1.23Y = 1.23(53.6) = 65.9

The quantities of X and Y which maximize utility (Uxy) = (65.9, 53.6)

 

C.). Maximum Utility:

 

U= X1/3Y2/3

U = 65.91/353.62/3

U = 57.4

Maximum Utility that the consumer can derive from x and y = 57.4